An oriental rug is 3 feet longer than it is wide. If the diagonal of the rug is 12 feet, find its dimensions to the nearest tenth of a foot.
Answer by Alex.33(110)

width=x
Pythagoreans theorem: x^2+(x+3)^2=12^2
x=6.9 ft, approximately
x+3=9.9 ft, approximately
dimensions:6.9x9.9
Done.